Coffee worth $3.75 a pound was mixed with coffee worth $4.35 a pound to produce a blend worth $4.11 a pound. How much of each kind of coffee was used to produce 40 pounds of blended coffee?
Answer by TimothyLamb(4379)

---
x = pounds of 3.75 coffee
y = pounds of 4.35 coffee
---
3.75x + 4.35y = 4.11*40
x + y = 40
---
put the system of linear equations into standard form:
---
3.75x + 4.35y = 164.4
x + y = 40
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = pounds of 3.75 coffee = 16
y = pounds of 4.35 coffee = 24
---
system matrix
3.75 4.35
1 1
---
inverse of system matrix
-1.6666667 7.25
1.6666667 -6.25
---
determinant of system matrix = -0.6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php